# Regularization methods for solving hierarchical variational inequalities with complexity guarantees

### v1

@@ 1 Introduction
Bi-level optimization problems consist of two nested optimization formulations, referred to as the inner and the outer problem. Solutions to the inner problem determine the feasible set over which a minimizer of an outer objective function is searched for. This class of problems is an extremely active area of research recently, driven by important practical applications in engineering, economics and machine learning. For recent comprehensive reviews, see [dempe2020bilevel, liu2021investigating]. A particularly appealing class of bi-level optimization problems are those in which only a single decision variable is involved at both layers of the problem. These special cases are often called simple bi-level optimization problems [Dempe:2021aa]. In this special instance, the problem is given by the convex constrained minimization problem
@@
Hence, if we set the operator $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}}$, Problem (P) specializes to a hierarchical variational inequality problem given by
@@
For recent studies of such problems in finite-dimensional euclidean spaces, we refer to [alves_inertial_2025] and [samadi_improved_2025].
@@
If $g$ is convex and Fréchet differentiable, this becomes a hierarchical variational inequality with $\mathsf{G}=\nabla g$, hence a particular instance of Problem (P). This viewpoint covers a broad range of models encountered in, e.g., engineering, economics, inverse problems and control (see [hintermuller2011first, hintermuller2015bi-level, hintermuller2014several, outrata2000generalized, de2023bi-level]).
@@ 1.1 Contributions
Numerical schemes with complexity statements for resolving Problem (P) have been recently published in the papers [Kaushik:2021aa, samadi_improved_2025] and [alves_inertial_2025]. [Kaushik:2021aa] extend the simple bi-level setting to a lower level problem defined in terms of a variational inequality while keeping the upper level problem as a convex optimization problem. This is a special instance of Problem (P), which is related to mathematical programming with equilibrium constraints [Luo_Pang_Ralph_1996], and to the equilibrium selection problem in game theory [Benenati:2022aa, Benenati:2023aa]. The recent papers [samadi_improved_2025, alves_inertial_2025] go one step further and study the special instance of Problem (P) in which the upper and lower level problem are defined in terms of a variational inequality problem.
In this paper we generalize the aforementioned existing results by considering a more general class of hierarchical equilibrium problems admitting the formulation (P). We prove rates on both levels of the equilibrium problem via a very flexible numerical scheme, which can easily be modulated with respect to problem structure (specifically of the lower-level inclusion problem). Our main contributions are threefold:
@@
  • Monotone-inclusion lower level and modular oracle. Rather than considering a variational inequality at the lower level, we allow the feasible set to be defined as the solution set of an abstract monotone inclusion. This formulation allows to draw from the rich theory of monotone operators and enables a very general algorithmic design based on a monotone-inclusion oracle obtained through a fixed-point encoding, which can be tailored to the structure of the operators involved.
  • Quantitative guarantees via gap functions. While much of the existing literature focuses on weak convergence results without explicit rate statements, we introduce gap functions for both the lower- and upper-level problems and establish quantitative convergence guarantees for these measures. Our results are formulated within a specific geometric setting in which we assume that the feasibility gap function is assumed to satisfy an Hölderian error bound condition such as [Cabot2005] in the context of hierarchical optimization and more recently [samadi_improved_2025] and [boct2025accelerating].
We achieve these results via the careful design of an algorithmic template drawing inspiration from classical diagonal methods involving iterative Tikhonov and proximal regularization via anchor terms. By constructing a family of auxiliary combined equilibrium problems, our method constructs a sequence in an inner loop method, which is designed to track a sequence of temporal solutions. This procedure is essentially an inertial Krasnoselskii-Mann iteration and thus only relies on the availability of a ready-to-compute fixed-point encoding map for the lower-level solution set $\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$. An outer-loop procedure updates the Tikhonov parameter and the anchor term to steer the trajectory of the numerical algorithm towards a solution of the entire Problem (P).
Diagonal schemes for solving convex programs are a classical topic [alart1991penalization]. The idea of extending this approach to hierarchical equilibrium problems is taken from [Facchinei:2014aa] and [Lampariello:2020aa]. In particular, [Lampariello:2020aa] also proposes a kind of double-loop scheme with a forward-backward iteration acting in the inner loop. In this paper, we consider a more general class of splitting problems, with explicit rate guarantees at both levels of the problem.
Specifically, in the case where $\mathsf{G}$ is merely monotone (opposed to strongly monotone), we obtain weak convergence of the averaged sequence of iterates $(\overline{w}_{N})$, with rates of the form, for arbitrary $b\in(0,1)$,
@@
for suitably defined optimality and feasibility gaps (see Section 3.2 for precise definition). In this setting, $\tilde{\mathcal{O}}$ disregards logarithmic factors, which are natural consequences of the double-loop structure of our algorithm. Ignoring these logarithmic factors, which previous works do not obtain due to single-loop structures, our work is on par with established benchmarks:
  • In [Kaushik:2021aa, Corollary 3.5], the authors study optimization problems with Cartesian variational inequality constraints, where the objective is convex and the constraint set is a Cartesian product associated with a monotone mapping. Their algorithm resembles a forward-backward method, cast into a single-loop scheme. For the deterministic variant of their method, they obtain rates of $\mathcal{O}(1/N^{0.5-b})$ for the optimality gap and $\mathcal{O}(1/N^{b})$ for the feasibility gap, for $b\in(0,0.5)$. Therefore, ignoring the logarithmic factors, the rates achieved in the present work are more favorable.
@@
$\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N}\right),\qquad\text{and}\qquad 0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N}\right).$
These rates are comparable to those in [samadi_improved_2025, Theorem 4.6], where a regularized extragradient method is analyzed. To the best of our knowledge, no other works address hierarchical programming problems in the strongly monotone setting in a manner that allows a direct comparison.
@@ 1.2 Notation and Definitions
Let $\mathcal{H}$ be a real Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and corresponding norm $\|{\cdot}\|$. A set-valued operator $\mathsf{A}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is monotone if $\langle u-v,b-a\rangle\geq 0$ for all $(u,b),(v,a)\in\operatorname{Gr}(\mathsf{A})\triangleq\{(v,a)\in\mathcal{H}\times\mathcal{H}|a\in\mathsf{A}(v)\}$. The operator $\mathsf{A}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is maximally monotone if it is monotone and there exists no other monotone operator whose graph properly contains $\operatorname{Gr}(\mathsf{A})$. The resolvent of an operator $\mathsf{A}$ is defined as $\operatorname{\mathsf{J}}_{\mathsf{A}}=(\operatorname{Id}+\mathsf{A})^{-1}$. If $\mathsf{A}$ is maximally monotone, then the resolvent is a nonexpansive operator and thus single-valued. A mapping $\mathsf{T}\colon\mathcal{H}\to\mathcal{H}$ is quasi-contractive if $\|Tx-p\|\leq q\|x-p\|$ for all $p\in\text{Fix}(T)$ and some $q\in(0,1)$. We refer to [BauCom16] for more details on operator theory.
For two sequences $(a_{n})\subset\mathbb{R}$ and $(b_{n})\subset\mathbb{R}$, we say that $a_{n}=O(b_{n})$ if there exists a constant $C>0$ such that $a_{n}\leq Cb_{n}$ for $n$ sufficiently large. We say that $a_{n}=o(b_{n})$ if $\lim_{n\to\infty}{b_{n}}/{a_{n}}=0$. Finally, we say that $a_{n}=\Theta(b_{n})$ if there exists a constant $C>0$ such that $\lim_{n\to\infty}{b_{n}}/{a_{n}}=C$.
@@ 2 Outline of the Algorithm
Remark 2.2.
We emphasize that our assumption includes the case $\mu=0$, in which $\mathsf{G}$ is only Lipschitz and monotone.
@@
Our numerical scheme consists of two main procedures. The inner procedure is an inertial Krasnoselskii-Mann (KM) iteration, whose last iterate is designed to lie in a sufficiently small ball around the temporal solution $\overline{u}_{\beta}(w)$, whose radius is part of the parameters of the procedure. We use this last iterate as a new anchor point, and restart the KM iteration with fresh parameters. To this end, we assume access to a family of parameterized fixed-point mappings $\mathsf{T}_{k}\equiv\mathsf{T}^{(w,\beta)}_{k}\colon\mathcal{H}\to\mathcal{H}$, which satisfy the following condition.
@@
Concrete examples of such operators are provided in Section 4. Generally, the operators may vary over the inner loop as a result of a varying step-size in the selected method.
@@
The proximal regularization parameter $\alpha>0$ is treated in our approach as an external input to the algorithm. Hence, its concrete numerical value used in the computation affects the central funnel $\overline{u}_{\beta}(w)$, as well as the fixed-point encoding map $\mathsf{T}^{(w,\beta)}_{k}$ and eventually also the fixed-point transportation map $\mathsf{Z}^{(w,\beta)}$. However, since it is a fixed parameter, we simplify the notation and omit an explicit dependence from the involved mappings and operators on the value of $\alpha$.
@@ 2.1 Inner Loop
The inner loop of our tracking algorithm employs the procedure $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$, see Algorithm 1, which is an inertial Krasnoselskii-Mann iteration with user-provided parameters embodied in terms of the initial point $v\in\mathcal{H}$, fixed-point encoding map $\mathsf{T}^{(w,\beta)}_{\bullet}$ satisfying Assumption 2, as well as momentum parameters $(\tau_{k})_{k}$ and relaxation parameters $(\theta_{k})_{k}$. This fixed-point iteration is designed to return an approximation of the central funnel fixed-point $\overline{u}_{\beta}(w)$ via the fixed-point transportation map $\mathsf{Z}^{(w,\beta)}$.
Algorithm 1 Function $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$
$v_{1}=v_{0}:=v$
@@
$\mathbb{K}(\varepsilon)\triangleq\inf\{k\geq 1|\;\|{v_{k+1}-z_{k}}\|\leq\varepsilon\}.$ (Stop)
This stopping criterion is motivated by the following approximation result.
@@
Let $(\tau_{k})_{k}\subset[0,1],(\theta_{k})_{k}\subset[0,\overline{\theta}]$ for $\overline{\theta}\in(0,1)$, $\varepsilon>0$, and let $(\mathsf{T}_{k})_{k}$ be a sequence of operators satisfying Assumption 2. Denote by $p$ the common fixed-point of $(\mathsf{T}_{k})_{k}$. If $\mathbb{K}(\varepsilon)$ is finite and $v_{\mathbb{K}(\varepsilon)+1}$ is the result of Algorithm 1, then
@@
Given $\varepsilon>0$, denote $\mathbb{K}=\mathbb{K}(\varepsilon)$. By definition of $v_{k+1}$ and quasi-contractiveness of $\mathsf{T}_{k}$, we get
@@
$\leq\frac{\varepsilon}{\theta_{\mathbb{K}}}\left({1-\theta_{\mathbb{K}}}+\frac{\overline{q}}{1-\overline{q}}\right)\leq\frac{\varepsilon}{\overline{\theta}}\left({1-\overline{\theta}}+\frac{\overline{q}}{1-\overline{q}}\right),$
as wanted. $\blacksquare$
Lemma 2.1 has important consequences for the tracking properties of the anchor points $w$ which we recursively generate with the restarting procedure explained next. To give an outlook, consider the sequence $(v_{k})_{k}$ produced by the function $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$, and let $w^{+}\triangleq\mathsf{Z}^{(w,\beta)}(v_{\mathbb{K}(\varepsilon)+1})$. Lemma 2.1 guarantees that this point satisfies the tracking guarantee
@@
by nonexpansiveness of $\mathsf{Z}^{(w,\beta)}$. The upper bound is thus of order $O(\varepsilon)$, and thus controlled by the user-defined targeted precision.

### v2

@@ 1 Introduction
Bi-level optimization problems consist of two nested optimization formulations, referred to as the inner and the outer problem. Solutions to the inner problem determine the feasible set over which a minimizer of an outer objective function is searched for. This class of problems is an extremely active area of research recently, driven by important practical applications in engineering, economics and machine learning. Comprehensive reviews can be found in [dempe2020bilevel, liu2021investigating]. A particularly appealing class of bi-level optimization problems are those in which only a single decision variable is involved at both layers of the problem. These special cases are often called simple bi-level optimization problems [Dempe:2021aa]. In this special instance, the problem is given by the convex constrained minimization problem
@@
Hence, by setting $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}}$, Problem (P) specializes to a hierarchical variational inequality problem
@@
For recent studies of such problem formulations in finite-dimensional Euclidean spaces, we refer to [alves_inertial_2025] and [samadi_improved_2025].
@@
If $g$ is convex and Fréchet differentiable, this becomes a hierarchical variational inequality with $\mathsf{G}=\nabla g$, hence a particular instance of Problem (P). This viewpoint covers a broad range of models encountered in, e.g., engineering, economics, inverse problems, and control (see [hintermuller2011first, hintermuller2015bilevel, hintermuller2014several, outrata2000generalized, de2023bilevel]).
@@ 1.1 Contributions
Numerical schemes with complexity statements for resolving Problem (P) have been recently published in the papers [Kaushik:2021aa, samadi_improved_2025] and [alves_inertial_2025]. [Kaushik:2021aa] extend the simple bi-level setting to a lower level problem defined in terms of a variational inequality, while keeping the upper level problem as a convex optimization problem. This is a particular case of Problem (P), which is related to mathematical programming with equilibrium constraints [Luo_Pang_Ralph_1996], and to the equilibrium selection problem in game theory [Benenati:2022aa, Benenati:2023aa]. The recent papers [samadi_improved_2025, alves_inertial_2025] go one step further and study a special instance of Problem (P) in which the upper- and lower-level problems are defined in terms of a variational inequality.
In this paper we generalize the aforementioned existing results, by considering a significantly larger class of hierarchical equilibrium problems, admitting the formulation (P). We prove rates on both levels of the equilibrium problem via a very flexible numerical scheme, which can easily be modulated to prevailing problem structure (specifically of the lower-level inclusion problem). Our main contributions are threefold:
@@
  • Monotone-inclusion lower level and modular oracle. Rather than considering a variational inequality at the lower level, we allow the feasible set to be defined as the solution set of an abstract monotone inclusion problem. Within this general framework, we adopt a very general algorithmic design based on a monotone-inclusion oracle obtained through a fixed-point encoding, which can be tailored to the structure of the operators involved. As a result, our numerical scheme and our analysis is applicable to entire families of relevant splitting methods, and not just a single one.
  • Quantitative guarantees via gap functions. While much of the existing literature focuses on weak convergence results without explicit rate statements, we introduce gap functions for both the lower- and upper-level problems and establish quantitative convergence guarantees for these measures. Our results are formulated within a specific geometric setting in which we assume that the gap function associated with the lower-level problem satisfies a Hölderian error bound condition. Such settings have already been considered in [Cabot2005] in the context of hierarchical optimization, and more recently in [samadi_improved_2025] and [boct2025accelerating].
We achieve these results via the careful design of an algorithmic template drawing inspiration from classical diagonal methods, involving iterative Tikhonov and proximal regularization via anchor terms. By constructing a family of auxiliary combined equilibrium problems, our method constructs a sequence in an inner loop method, which is designed to track a sequence of temporal solutions. This procedure is essentially an inertial Krasnoselskii-Mann (KM) iteration and thus only relies on the availability of a ready-to-compute fixed-point encoding map for the lower-level solution set $\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$. An outer-loop restarting procedure updates the Tikhonov parameter and the anchor term to steer the trajectory of the numerical algorithm towards a solution of (P).
Diagonal schemes for solving convex programs are rather classical (see e.g. [alart1991penalization]). The idea of extending this approach to hierarchical equilibrium problems is taken from [Facchinei:2014aa] and [Lampariello:2020aa]. In particular, [Lampariello:2020aa] also proposes a kind of double-loop scheme with a forward-backward algorithm acting in the inner loop. In this paper, we consider a more general class of splitting problems, with explicit rate guarantees at both levels of the problem. Specifically, in the case where $\mathsf{G}$ is merely monotone (as opposed to strongly monotone), we obtain weak convergence of the averaged sequence of iterates $(\overline{w}_{n})_{n\in\mathbb{N}}$, with rates of the form, for arbitrary $b\in(0,1)$,
@@
for suitably defined optimality and feasibility gaps (see Section 3.2 for precise definition). In this setting, the integer $n$ measures the number of restarts, and $\tilde{\mathcal{O}}$ disregards logarithmic factors, which are consequences of the double-loop structure of our algorithm. Besides these logarithmic factors, our work is on par with published single-loop algorithms:
  • In [Kaushik:2021aa, Corollary 3.5], the authors study optimization problems with Cartesian variational inequality constraints, where the objective is convex and the constraint set is a Cartesian product associated with a monotone mapping. Their algorithm resembles a forward-backward method, cast into a single-loop scheme. For the deterministic variant of their method, they obtain rates of $\mathcal{O}(1/N^{0.5-b})$ for the optimality gap and $\mathcal{O}(1/N^{b})$ for the feasibility gap, for $b\in(0,0.5)$, $N$ being the iteration count. Therefore, ignoring logarithmic factors, the rates achieved in the our work are more favorable.
@@
$\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N}\right),\qquad\text{and}\qquad 0\leq\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})\leq\tilde{\mathcal{O}}\left(\frac{1}{N}\right),$
where, again, $N$ measures the number of restarts. These rates are comparable to those in [samadi_improved_2025, Theorem 4.6], where a regularized extragradient method is analyzed.
@@ 1.2 Notation and Definitions
We adopt standard notation and terminology from variational analysis, as in [BauCom16]. Let $\mathcal{H}$ be a real Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and corresponding norm $\|{\cdot}\|$. A set-valued operator $\mathsf{A}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is $\mu$-monotone if $\mu\geq 0$ and
$\langle b-a,u-v\rangle\geq\mu\|{u-v}\|^{2}\text{ for all }(u,b),(v,a)\in\operatorname{Gr}(\mathsf{A})\triangleq\{(v,a)\in\mathcal{H}\times\mathcal{H}|a\in\mathsf{A}(v)\}.$
If $\mu=0$, we will say $\mathsf{A}$ is merely monotone. The operator $\mathsf{A}\colon\mathcal{H}\to 2^{\mathcal{H}}$ is maximally monotone if it is monotone and there exists no other monotone operator whose graph properly contains $\operatorname{Gr}(\mathsf{A})$. The domain of an operator $\mathsf{A}$ is $\operatorname{dom}(\mathsf{A})=\{x\in\mathcal{H}|\mathsf{A}(x)\neq\varnothing\}$. A mapping $\mathsf{T}\colon\mathcal{H}\to\mathcal{H}$ is quasi-contractive if $\|Tx-p\|\leq q\|x-p\|$ for all $p\in\textnormal{Fix}(T)$ and some $q\in(0,1)$, where $\textnormal{Fix}(\mathsf{T})\triangleq\{x\in\mathcal{H}|x=\mathsf{T}(x)\}$ is the set of fixed points of $\mathsf{T}$. If the same inequality holds for $q=1$, then $\mathsf{T}$ is non-expansive.
The resolvent of an operator $\mathsf{A}$ is defined as $\operatorname{\mathsf{J}}_{\mathsf{A}}\triangleq(\operatorname{Id}+\mathsf{A})^{-1}$. By Minty’s theorem, if $\mathsf{A}$ is maximally monotone, then the resolvent is a nonexpansive operator and thus single-valued. The normal cone mapping to a closed convex and non-empty set $\mathcal{K}\subset\mathcal{H}$ is defined as $\operatorname{\mathsf{NC}}_{\mathcal{K}}(x)=\{u\in\mathcal{H}|\langle u,y-x\rangle\leq 0\;\forall y\in\mathcal{K}\}$ for $x\in\mathcal{K}$ and $\operatorname{\mathsf{NC}}_{\mathcal{K}}(x)=\varnothing$ otherwise. Under convexity, $\operatorname{\mathsf{NC}}_{\mathcal{K}}$ is maximally monotone, and its resolvent coincides with the orthogonal projector $\Pi_{\mathcal{K}}(x)=\operatorname*{argmin}\{\|{y-x}\|:y\in\mathcal{K}\}$.
For two sequences $(a_{n})\subset\mathbb{R}$ and $(b_{n})\subset\mathbb{R}$, we say that $a_{n}={\mathcal{O}}(b_{n})$ if there exists a constant $C>0$ such that $a_{n}\leq Cb_{n}$ for $n$ sufficiently large. We say that $a_{n}=o(b_{n})$ if $\lim_{n\to\infty}{b_{n}}/{a_{n}}=0$. Finally, we say that $a_{n}=\Theta(b_{n})$ if there exists a constant $C>0$ such that $\lim_{n\to\infty}{b_{n}}/{a_{n}}=C$.
@@ 2 Outline of the Algorithm
Our numerical scheme consists of two main procedures. The inner procedure is an inertial KM iteration, whose last iterate is designed to lie in a sufficiently small ball around the temporal solution $\overline{u}_{\beta}(w)$, with iteratively defined radius. We subsequently use this last iterate as a new anchor point, and restart the KM iteration with fresh parameters. To this end, we assume to have access to a family of parameterized fixed-point mappings $\mathsf{T}_{k}\equiv\mathsf{T}^{(w,\beta)}_{k}\colon\mathcal{H}\to\mathcal{H}$, satisfying the following condition.
@@
Concrete examples of such fixed-point encoding mappings are provided in Section 4.
@@
The proximal regularization parameter $\alpha>0$ is treated in our approach as an external input to the algorithm. Hence, its concrete numerical value used in the computation affects the position of the solution $\overline{u}_{\beta}(w)$, the shape of the fixed-point encoding map $\mathsf{T}^{(w,\beta)}_{k}$, and eventually also the fixed-point transportation map $\mathsf{Z}^{(w,\beta)}$. However, since it is a fixed parameter, we simplify the notation and omit an explicit dependence from the involved mappings and operators on the value of $\alpha$.
@@ 2.1 Inner Loop
The inner loop of our tracking algorithm employs the procedure $\mathtt{KM}(v,(\mathsf{T}^{(w,\beta)}_{k})_{k},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$, see Algorithm 1, which is an inertial Krasnoselskii-Mann iteration with user-provided parameters embodied in terms of the initial point $v\in\mathcal{H}$, fixed-point encoding mappings $(\mathsf{T}^{(w,\beta)}_{k})_{k}$ satisfying Assumption 2, as well as momentum parameters $(\tau_{k})_{k}$ and relaxation parameters $(\theta_{k})_{k}$. This fixed-point iteration is designed to return an approximation of the central funnel fixed-point $\overline{u}_{\beta}(w)$ via the fixed-point transportation map $\mathsf{Z}^{(w,\beta)}$.
Algorithm 1 Function $\mathtt{KM}(v,(\mathsf{T}^{(w,\beta)}_{k})_{k},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$
$v_{1}=v_{0}\triangleq v$
@@
$\mathbb{K}(\varepsilon)\triangleq\inf\{k\geq 1|\;\|{v_{k+1}-z_{k}}\|\leq\varepsilon\}.$ (2.1)
This stopping criterion is motivated by the following approximation result, formulated in terms of general fixed-point encodings obeying the imposed assumptions.
@@
Let $(\tau_{k})_{k}\subset[0,1],(\theta_{k})_{k}\subset[0,\overline{\theta}]$ for $\overline{\theta}\in(0,1)$, $\varepsilon>0$, and let $(\mathsf{T}_{k})_{k}$ be a sequence of operators satisfying Assumption 2. Denote by $p$ the common fixed-point of $(\mathsf{T}_{k})_{k}$. If the stopping time $\mathbb{K}(\varepsilon)$ defined in (2.1) is finite and $v_{\mathbb{K}(\varepsilon)+1}$ is the last iterate produced by Algorithm 1, then
@@
Given $\varepsilon>0$, let $\mathbb{K}=\mathbb{K}(\varepsilon)$. By definition of $v_{k+1}$ and quasi-contractiveness of $\mathsf{T}_{k}$, we get
@@
$\leq\frac{\varepsilon}{\theta_{\mathbb{K}}}\left({1-\theta_{\mathbb{K}}}+\frac{\overline{q}}{1-\overline{q}}\right)\leq\frac{\varepsilon}{\overline{\theta}}\left({1-\overline{\theta}}+\frac{\overline{q}}{1-\overline{q}}\right).$
$\blacksquare$
Lemma 2.1 has important consequences for the tracking properties of the anchor points $w$ which we recursively generate with the restarting procedure explained next. To give an outlook, consider the sequence $(v_{k})_{k}$ produced by the function $\mathtt{KM}(v,(\mathsf{T}^{(w,\beta)}_{k})_{k},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$, and let $w^{+}\triangleq\mathsf{Z}^{(w,\beta)}(v_{\mathbb{K}(\varepsilon)+1})$. Lemma 2.1 guarantees that this point satisfies the tracking guarantee
@@
by nonexpansiveness of $\mathsf{Z}^{(w,\beta)}$. The upper bound is of order $\mathcal{O}(\varepsilon)$, and thus controlled by the user-defined targeted precision.